OpenStudy (anonymous):

5 6 - = ------------ 3 x+2 SOLVE FOR X. NEED THE STEPS

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

\[(5/3) = (6/(x+2))\] multiply both side by (x+2, then multiply both sides by 3 \[(X+2)*5 = 6*3\] now divide both sides by 5 \[(X+2) = 18/5\] subtract two from both sides \[(x+2)-2 = [(18/5)-(10/5)]\] now you have \[x=8/5\]

OpenStudy (anonymous):

Move the RHS to the LHS.\[\frac{5}{3}-\frac{6}{x+2}=0 \]\[\frac{a}{b}-\frac{c}{d}=\frac{a d-b c}{b d}\]Set ad -bc to zero and solve for x.\[5(x+2)-3*6=0 \]\[5 x-8=0 \]\[x=\frac{8}{5} \]Same as deezil_wood 's answer.